I need to solve for y to graph. The equation is x-2y=10. How do I do that?
Answer by JBarnum(2146)

x-2y=10
-x -x
-2y=10-x
(-2y)/-2=(10-x)/-2
y=(x-10)/2 or y=(x/2)-5